# appendixReferences

### v3

@@ Abstract
We give an approach for characterizing interference by lower bounding the number of units whose outcome depends on certain groups of treated individuals, such as depending on the treatment of others, or others who are at least a certain distance away. The approach is applicable to randomized experiments with binary-valued outcomes. Asymptotically conservative point estimates and one-sided confidence intervals may be constructed with no assumptions beyond the known randomization design, allowing the approach to be used when interference is poorly understood, or when an observed network might only be a crude proxy for the underlying social mechanisms. Point estimates are equal to Hájek-weighted comparisons of units with differing levels of treatment exposure. Empirically, we find that the width of our interval estimates is competitive with (and often smaller than) those of the EATE, an assumption-lean treatment effect, suggesting that the proposed estimands may be intrinsically easier to estimate than treatment effects.
@@ 1 Introduction
There is growing interest in the analysis of randomized experiments where participants may affect each other (“interference between units”), for example by transmission of disease, sharing of information, peer influence, or economic competition. Such mechanisms often play fundamental roles in our understanding of social behaviors, and their empirical characterization may be of scientific interest. In this work, we give a general approach for characterizing the prevalence of treatment effects and interference, which can be used to evaluate whether interference-causing mechanisms are widely prevalent in an experimental setting. The approach is applicable to experiments with binary-valued outcomes. It can be used to explore questions such as
@@
    How many units are affected by any treatment (including their own)?
@@
    How many units are affected by the treatment of others? of distant others?
  3. 3.
    How many units are affected by the treatment of others, when their own treatment satisfies some condition?
For each of these questions, our approach gives conservative point estimates and one-sided confidence intervals, both of which lower bound the true value. Under reasonable experiment designs, our point estimates will converge to a lower bound on the true value, while the one-sided interval will have coverage probability converging to at least the desired level. These consistency and coverage properties are agnostic, in that they require no assumptions or restrictions on the nature of the interference, and rely only on the known experiment design. As a result, our estimates will remain valid even if an observed social network is only a crude proxy for the actual social mechanisms.
Existing work in this area produced very loose (though valid) lower bounds (Choi, , 2023). Our approach is significantly tighter, and may be more practical as a result. After Hájek normalization, our point estimates are equal to Hájek-weighted contrasts, such as comparisons of treated versus untreated, or comparisons of different levels of indirect exposure. Under stronger assumptions, contrasts arise naturally as estimates of treatment effects. Our results indicate that without assumptions, these contrasts may sometimes be re-interpreted more weakly as lower bounds on the number of units who are affected by the treatments in a particular manner, albeit with new confidence intervals that use a novel bound for the estimator variance. We find empirically that the uncertainty of our estimates (i.e., the interval widths) is competitive with, and often better than, that of the EATE, an assumption-lean treatment effect.
Our approach does not replace estimation of treatment effects, but may serve as a useful complement. In the presence of interference, estimates of treatment effects may come with considerable caveats or limitations – for example by requiring strong assumptions, or by estimating marginal effects rather than total ones. As a result, the usage of these estimates may require caution, judgement, and a reliance on prior beliefs. Agnostic information about the underlying system may help by informing and challenging these beliefs. If the decision maker is confronted by the results of an agnostic analysis, even one that does not directly answer the policy question of interest, then their beliefs may require recalibration in light of the findings, which in turn may affect downstream decisions. On the other hand, an assumption-laden analysis may carry less weight if caution is required to interpret its findings.
@@ 2 Related Work
Our work is closely related to hypothesis testing in the presence of interference, which has been studied in Aronow, (2012); Athey et al., (2018); Pouget-Abadie et al., (2019); Basse et al., (2019); Han et al., (2022); Puelz et al., (2022); Hoshino and Yanagi, (2023), for testing of null effects and also of hypotheses regarding interference structures, such as the null of no interference between units. In these settings, p-values generally cannot be inverted into confidence intervals for a treatment effect unless additional assumptions are added, and the connection between testing and estimation under interference is not fully understood. Our work complements this literature, as we invert tests of interference specifications to construct one-sided intervals without additional assumptions – however not for a treatment effect, but rather for the number of units who violate the specification.
The estimation of treatment effects in the presence of interference is an active area of research (Sävje et al., , 2021; Li et al., , 2019; Hu et al., , 2021; Li and Wager, , 2022; Leung, , 2022; Ogburn et al., , 2022; Harshaw et al., , 2022; Park and Kang, , 2023; Gao and Ding, , 2023). In many of these works, the target parameter describes the effect of a marginal departure from the experiment design. For such estimands, consistent estimates can generally be constructed in assumption-lean regimes, although additional assumptions are needed for coverage of interval estimates. Harshaw et al., (2021) focuses on the problem of variance estimation. Estimation of total effects (as opposed to marginal effects) assuming low order interactions between unit treatments is considered in Cortez et al., (2022) under staggered rollout, and in Cortez-Rodriguez et al., (2023) for known exposure neighborhoods. Blake and Coey, (2014) describe a particular form of interference arising in e-commerce settings, studied in Farias et al., (2022); Dhaouadi et al., (2023); Johari et al., (2024). Banerjee et al., (2024) discusses an empirical campaign for information sharing in which interference may be more subtle and complex than is captured by models of local exposure such as infection.
@@ 3 Methodology (basic case)
Our approach can be adapted to a variety of estimands that characterize interference, including the questions that were presented in the introduction. In this section, we give a basic example, in which the estimand is the number of units who are affected by any treatment. This quantity, which we denote by $\tau^{\textup{basic}}$, includes units who are affected by their own treatment as well as those who are affected by the treatment of others. Sections 3.1 through 3.6 respectively present the generative model, estimand, point estimation, Hájek normalization, inference, and computation. In Section 4, the approach is then generalized to cover a wider range of target parameters.
3.1 Preliminaries
@@
Given an observed network, we will use $\eta_{i}$ to denote the set consisting of unit $i$ and $i$’s “close” neighbors, and $\eta_{i}^{(2)}$ to denote “non-close” neighbors of unit $i$. For example, $\eta_{i}$ might denote unit $i$ and those units whose shortest path distance to $i$ is 1, while $\eta_{i}^{(2)}$ might denote those whose shortest path distance is 2. We will let $W_{i}$ denote the thresholded number of treated close neighbors for unit $i$,
$W_{i}=1\{\textup{if at least $t_{i}$ units in $\eta_{i}$, excluding unit $i$, are treated}\},$ (2)
and $W_{i}^{(2)}$ will denote the thresholded number of treated non-close neighbors,
$W_{i}^{(2)}=1\{\textup{ if at least $t_{i}^{(2)}$ units in $\eta_{i}^{(2)}$ are treated}\}$ (3)
@@
Given integer $n>0$, we use $[n]$ to denote the set $\{1,\ldots,n\}$. Given a set $S\subseteq[N]$, we use $X_{S}$ denote the subvector of $X$ corresponding to the members of $S$.
@@ 3.2 Estimand
Let $\phi^{*}=(\phi_{1}^{*},\ldots,\phi_{N}^{*})$ indicate whether each unit is unaffected by treatment,
$\phi_{i}^{*}=\begin{cases}1&\textup{if $f_{i}(X)$ is constant in $X$}\\ 0&\textup{otherwise}\end{cases}$
Equivalently, we may say that $Y_{i}$ is deterministic if $\phi_{i}^{*}=1$. Let $\Phi$ denote the number of units unaffected by treatment,
$\Phi=\sum_{i=1}^{N}\phi_{i}^{*}$
and let $\tau^{\textup{basic}}$ denote the complement of $\Phi$,
$\tau^{\textup{basic}}=N-\Phi$
which equals the total number of units who are affected by $X$. The target parameter of interest is $\tau^{\textup{basic}}$.
@@ 3.3 Point Estimation
To estimate $\tau^{\textup{basic}}$, we first consider two Horvitz-Thompson estimators for $\Phi$, which are not computable as they require knowledge of $\phi^{*}$:
$\hat{\Phi}_{1}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=Y_{i}\}}{{P}(X_{i}=Y_{i})}$ $\hat{\Phi}_{2}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}\neq Y_{i}\}}{{P}(X_{i}\neq Y_{i})}$ (4)
To give a single formula for both estimators we may define $Y_{ki}$ for $k\in[2]$ and $i\in[N]$ as the estimator-specific outcome
$Y_{ki}=\begin{cases}Y_{i}&\textup{ if $k=1$}\\ (1-Y_{i})&\textup{ if $k=2$}\end{cases}$
and rewrite (4) as
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=Y_{ki}\}}{P(X_{i}=Y_{ki})}$ (5)
To evaluate the probability $P(X_{i}=Y_{ki})$, we observe that if $\phi_{i}^{*}=1$, then $Y_{ki}$ is deterministic, implying that
$P(X_{i}=Y_{ki})=Y_{ki}\cdot P(X_{i}=1)+(1-Y_{ki})\cdot P(X_{i}=0),\qquad\textup{ if $\phi_{i}^{*}=1$.}$ (6)
It follows that (5) is an unbiased estimator of $\Phi$ provided that $P(X_{i}=0)$ and $P(X_{i}=1)$ are non-zero for all $i\in[N]$.
@@
Since only deterministic outcomes are involved, the estimators will often exhibit simple behavior. For example, if treatment is assigned by independent Bernoulli randomization, then $\hat{\Phi}_{k}$ is a sum of independent variables, regardless of any interference that may be present. Similarly, if the dependencies between the unit treatments are bounded, then $\hat{\Phi}_{k}$ is a sum of variables whose dependencies will be similarly bounded. For this reason, under a variety of designs we may expect $\hat{\Phi}_{k}$ to concentrate at its expectation and be asymptotically normal (while also being uncomputable as $\phi^{*}$ is unknown).
Let $\hat{\Phi}_{k}(\phi)$ denote $\hat{\Phi}_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$. To give a concrete expression, we may write
$\hat{\Phi}_{k}(\phi)=\sum_{i:\phi_{i}=1}\frac{1\{X_{i}=Y_{ki}\}}{Y_{ki}\cdot P(X_{i}=1)+(1-Y_{ki})\cdot P(X_{i}=0)}$ (7)
where the denominator term is due to application of (6).
To conservatively estimate $\tau^{\textup{basic}}=N-\Phi$, we will plug in for $\Phi$ the smaller of $\hat{\Phi}_{1}(\phi)$ and $\hat{\Phi}_{2}(\phi)$ after maximizing over $\phi\in\{0,1\}^{N}$, resulting in the estimator
$\hat{\tau}=N-\min\Big{(}\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}(\phi)\ ,\ \max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}(\phi)\Big{)}$ (8)
The main properties of $\hat{\tau}$ are the following:
  1. 1.
    If $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ are both consistent for $\Phi$, then $\min(\hat{\Phi}_{1},\hat{\Phi}_{2})$ will also be consistent for $\Phi$, and $\hat{\tau}$ will asymptotically lower bound $\tau^{\textup{basic}}$, as stated by Theorem 1 given below:
    Theorem 1.
    Let $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ be consistent for $\Phi$, so that
    $\hat{\Phi}_{k}=\Phi+o_{P}(N),\qquad k=1,2$ (9)
    Then $\hat{\tau}$ converges to a lower bound on $\tau^{\textup{basic}}$, so that
    $\hat{\tau}\leq\tau^{\textup{basic}}+o_{P}(N)$ (10)
  2. 2.
    Despite its conservative nature as a lower bound, $\hat{\tau}$ will take large values when the difference in average outcomes is large between treated and control. To illustrate, Figure 1 shows an experiment in which the control units have substantially higher outcomes than the treated. As can be seen, the large difference in outcomes forces the set $\{i:X_{i}=Y_{i}\}$ to be small in size, and equivalently |$|\{i:X_{i}=Y_{ki}\}|$ must be small for $k=1$. In turn, this forces $\hat{\Phi}_{1}(\phi)$ given by (7) to be small for all $\phi$, and hence forces $\hat{\tau}$ given by (8) to be large.
    In fact, under mild conditions $\hat{\tau}$ converges to the inverse propensity weighted (IPW) difference in average outcomes between treated and control, as stated by Theorem 2:
    Theorem 2.
    Let $X$ satisfy
    $\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}=N+o_{P}(N)\quad\textup{and}\quad\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}=N+o_{P}(N)$ (11)
@@
    so that the total weights of the treated and control converge to their expectations. Then $\hat{\tau}$ is asymptotically the IPW difference in means
    $\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\sum_{i=1}^{N}\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right|+o_{P}(N)$ (12)
Figure 1: Experiment in which control units have substantially higher outcomes than treated, resulting in a small number of units for which $X_{i}=Y_{i}$ (black outlined rectangles)

### v4

# Agnostic Characterization of Interference in Randomized Experiments

@@ Abstract
We give an approach for characterizing interference by lower bounding the number of units whose outcome depends on selected groups of treated individuals, such as depending on the treatment of others, or others who are at least a certain distance away. The approach is applicable to randomized experiments with binary-valued outcomes. Asymptotically conservative point estimates and one-sided confidence intervals may be constructed with no assumptions beyond the known randomization design, allowing the approach to be used when interference is poorly understood, or when an observed network might only be a crude proxy for the underlying social mechanisms. Point estimates are equal to Hájek-weighted comparisons of units with differing levels of treatment exposure. Empirically, we find that the width of our interval estimates is competitive with (and often smaller than) those of the EATE, an assumption-lean treatment effect, suggesting that the proposed estimands may be intrinsically easier to estimate than treatment effects.
@@ 1 Introduction
There is growing interest in the analysis of randomized experiments where participants may affect each other (“interference between units”), for example by transmission of disease, sharing of information, peer influence, or economic competition. Such mechanisms often play fundamental roles in our understanding of social behaviors, and their empirical characterization may be of scientific interest.
We believe that a better understanding of interference may not only be of scientific interest, but also of immense practical importance. One of the most well-known findings in the history of public health–that cholera could be transmitted through shared water utilities–was a particularly impactful characterization of interference (Freedman, , 2010, Ch. 3). Conversely, List, (2022) and List, (2023) discuss fields experiments at Uber in which the results were misinterpreted because interference was ignored, leading to decisions that harmed the company.111(List, , 2022, pp. 93-94) reports that giving a rider a promotional coupon increased their Uber usage in a field study, but the possibility of negative effects on other riders (by increasing their waiting time for rides) was ignored. The promotion was subsequently expanded to cover an entire city, leading to long wait times followed by substantial reduction in Uber usage. Similarly, List, (2023) reports that allowing a driver to receive tips increased their productivity in a field experiment, but the possibility of negative effects on the productivity of other drivers (who competed for the same pool of riders) was again ignored. Tipping was subsequently allowed for all drivers, but no gains in driver productivity were realized. The 2007 subprime mortgage crisis and subsequent “Great Recession” have been attributed in part to questionable modeling of causal dependencies between mortgage defaults, which led to catastrophic underestimation of credit risk (Salmon, , 2012). [^2] Interference between mortgages arises if housing prices can affect–and are affected by–the rate of mortgage default.
In this work, we give a general approach for characterizing whether interference-causing mechanisms are widely prevalent in an experimental setting. The approach is applicable to experiments with binary-valued outcomes. It can be used to explore questions such as
@@
    How many units are affected by (i.e., their outcome depends on) the treatment of others? of distant others?
@@
    How many units, when receiving a certain level of direct treatment, are affected by the treatment of others? For example, if a vaccine grants immunity then the disease status of a unit receiving the placebo (but not necessarily when receiving the vaccine), might depend on the vaccination status of others.
For these types of questions, our approach gives conservative point estimates and one-sided confidence intervals, both of which lower bound the true value. Under reasonable experiment designs, our point estimates will converge to a lower bound on the true value, while the one-sided interval will cover with probability converging to at least the desired level. These consistency and coverage properties are agnostic, in that they require no assumptions or restrictions on the nature of the interference, and rely only on the known experiment design. As a result, our estimates will remain valid lower bounds (but may become less tight) if an observed social network is only a crude proxy for the actual social mechanisms.
We will demonstrate our approach using an experiment of Cai et al., (2015), which studied whether crop insurance decisions of farmers could be affected by information received by their peers. Such experiments, in which the primary goal is to demonstrate the presence of interference, may be considered part of the collection of “shoe leather” efforts that are necessary to deeply understand a problem setting (Freedman, , 2010).333“Historically, ‘shoe-leather epidemiology’ is epitomized by intensive, door-to-door canvassing that wears out investigators’ shoes” (Freedman, , 2010, p. xiii). Without such efforts, treatment effects and policy recommendations may lead to questionable conclusions, such as in the abovementioned Uber examples where interference was mischaracterized. On the other hand, a more informed understanding of interference may lead to better assumptions and more credible policy recommendations. For the experiment of Cai et al., (2015), our results strengthen their findings, with confidence intervals that are both tighter and require fewer assumptions than those of the expected average treatment effect (EATE), an assumption-lean treatment effect whose study was motivated by interference (Sävje et al., , 2021).
@@ 2 Related Work
Our work is closely related to hypothesis testing in the presence of interference, which has been studied in Aronow, (2012); Athey et al., (2018); Pouget-Abadie et al., (2019); Basse et al., (2019); Han et al., (2022); Puelz et al., (2022); Hoshino and Yanagi, (2023), for testing of null effects and hypotheses regarding interference structures, such as the null of no interference between units. In these settings, p-values generally cannot be inverted into confidence intervals for a treatment effect unless additional assumptions are added, and more generally the connection between testing and estimation under interference is not fully understood. Our work complements this literature, as we invert tests of interference specifications to construct one-sided intervals without additional assumptions – however not for a treatment effect, but rather for the number of units who violate the specification.
In doing so, we improve on prior work in this area (Choi, , 2023, 2024) in multiple respects. Most importantly, our confidence intervals are much tigher than those of previous approaches, due to a combination of improved computation, better estimator design, and a novel approach for variance estimation. Our target estimand is a proper parameter, whereas in prior approaches it was defined to be a random quantity whose value depended on the treatment assignment. Finally, our approach generalizes to answer a wide variety of questions regarding the characterization of interference.
The estimation of treatment effects in the presence of interference is an active area of research (Rosenbaum, , 2007; Sävje et al., , 2021; Li et al., , 2019; Hu et al., , 2021; Li and Wager, , 2022; Leung, , 2022; Ogburn et al., , 2022; Harshaw et al., , 2022; Zhao and Ding, , 2022; Park and Kang, , 2023; Gao and Ding, , 2023). In many of these works, the target parameter describes the effect of a marginal departure from the experiment design. For such estimands, consistent estimates can generally be constructed in assumption-lean regimes, although additional assumptions are needed for coverage of interval estimates. Harshaw et al., (2021) focuses on the problem of variance estimation. For settings with interference, Kandiros et al., (2024) proposes a novel, more-tailored approach to experiment design. Estimation of total effects (as opposed to marginal effects) assuming low order interactions between unit treatments is considered in Cortez et al., (2022) under staggered rollout, and in Cortez-Rodriguez et al., (2023) for known exposure neighborhoods. Farias et al., (2022), Dhaouadi et al., (2023), and Johari et al., (2024) consider a particular form of interference (“Markovian”) arising in e-commerce settings. Banerjee et al., (2024) gives an empirical example of information sharing in which interference may be more subtle and complex than is captured by models of local exposure such as infection.
@@ 3 Methodology (basic case)
In this section, we present a basic example, in which the estimand is the number of units who are affected by (i.e., their outcome depends on) the assignment of treatment. This quantity, which we denote by $\tau^{\textup{basic}}$, includes units who are affected by their own treatment as well as those who are affected by the treatment of others.
Under certain experiment designs, $\tau^{\textup{basic}}$ may be used to characterize interference. In bipartite network designs, some units are not directly treated, and $\tau^{\textup{basic}}$ may be used to characterize whether these units are affected by the treatment of others (Zigler and Papadogeorgou, , 2020). In experiments involving the assignment of units to groups, or the formation of network ties, $\tau^{\textup{basic}}$ characterizes whether units are affected by their randomly assigned neighbors or group members (Basse et al., , 2024; Li et al., , 2019).
The remaining organization of the paper is as follows. Sections 3.1 through 3.7 respectively present the generative model, the estimand $\tau^{\textup{basic}}$, point estimation, usage of network information, Hájek normalization, inference, and computation. In Section 4, the approach is then generalized to cover a wider range of target parameters, such as the questions presented in the introduction. Section 5 considers asymptotic behavior, and Section 6 gives a simulation study and data example to illustrate usage. All propositions and theorems are proven in the supplemental materials.
3.1 Preliminaries and Oveview
@@
Given an observed network, we will use ${\mathcal{N}}_{i}$ to denote the set consisting of unit $i$ and $i$’s “close” associates, and ${\mathcal{N}}_{i}^{(2)}$ to denote “non-close” associates of unit $i$, as defined by the researcher. For example, ${\mathcal{N}}_{i}$ might denote unit $i$ and its direct neighbors on the network, while ${\mathcal{N}}_{i}^{(2)}$ might denote units who are not in $\mathcal{N}_{i}$ but have overlapping sets of neighbors with unit $i$. We will let $W_{i}$ denote the thresholded number of treated close associates for unit $i$,
[…]
and $W_{i}^{(2)}$ will denote the thresholded number of treated units in $\mathcal{N}_{i}^{(2)}$,
[…]
@@
Given integer $n>0$, we use $[n]$ to denote the set $\{1,\ldots,n\}$. Given a set $S\subseteq[N]$, we let $|S|$ denote its cardinality, and let $X_{S}$ denote the subvector of $X$ corresponding to the members of $S$.
Overview of Approach
To estimate lower bounds on our target estimand $\tau^{\textup{basic}}$, we will take the following approach:
  1. 1.
    Define the estimand $\tau^{\textup{basic}}$ in terms of an particular subset ${\mathcal{I}}\subseteq[N]$, whose identity is unknown
  2. 2.
    Propose idealized estimators $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ which will have good statistical properties, such as consistency and asymptotic normality, but require knowledge of $\mathcal{I}$}
  3. 3.
    Show that $\Delta$, the difference in average outcomes between treated and control (which can be computed without knowledge of $\mathcal{I}$}) converges to a lower bound for $\max(\hat{\tau}_{1},\hat{\tau}_{2})$, so that if $\tau_{1}$ and $\tau_{2}$ are both consistent for $\tau^{\textup{basic}}$, then $\Delta$ is an asymptotic lower bound.
  4. 4.
    Lower bound the confidence intervals induced by $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ and their variance estimates, by solving an integer program over the unknown subset $\mathcal{I}$}.
@@ 3.2 Estimand
Let $\tau^{\textup{basic}}$ denote the number of units who are affected by treatment, including both their own treatment or the treatment of others. We may write $\tau^{\textup{basic}}$ as the number of units whose outcome $f_{i}(X)$ is not constant over the possible treatment assignments $X\in\{0,1\}^{N}$.
$\tau^{\textup{basic}}=\sum_{i=1}^{N}1\{f_{i}(X)\textup{ is not constant in }X\}$
Define ${\mathcal{I}}\subset[N]$ as the subset of units who are unaffected by treatment and have constant outcome mappings,
${\mathcal{I}}=\{i:f_{i}(X)\textup{ is constant in }X\}$
so that $\tau^{\textup{basic}}=N-|{\mathcal{I}}|$.
@@ 3.3 Point estimation
Let $H_{i}$ denote the indicator of whether unit $i$’s treatment and outcome have the same binary value,
$H_{i}=1\{(X_{i},Y_{i})=(1,1)\textup{ or }(0,0)\},$ (4)
and let $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ denote sampling-based estimators of $\tau^{\textup{basic}}=N-|{\mathcal{I}}|$, in which the unknown cardinality of $\mathcal{I}$} is unbiasedly estimated by a probability-weighted (i.e., Horvitz-Thompson) sample:
$\hat{\tau}_{1}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{H_{i}=1\}}{P(H_{i}=1)}\qquad\textup{ and }\qquad\hat{\tau}_{2}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{H_{i}=0\}}{P(H_{i}=0)}$ (5)
@@
Because $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ involve only units in $\mathcal{I}$} whose outcomes are unaffected by treatment, they often will exhibit simple statistical behavior, even if strong interference exists between units who are not in $\mathcal{I}$}. For example, if treatment is assigned by independent Bernoulli randomization, then $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are sums of independent variables. [^4] If $i\in{\mathcal{I}}$, then $Y_{i}$ equals a constant $c_{i}$ for all $X\in\{0,1\}$. Thus it holds for example that $P(H_{i}=1,H_{j}=1)=P(X_{i}=c_{i},X_{j}=c_{j})=P(X_{i}=c_{i})P(X_{j}=c_{j})=P(H_{i}=1)P(H_{j}=1)$, if $i,j\in{\mathcal{I}}$ and $X_{i}\perp\!\!\!\!\perp X_{j}$. Similarly, if the dependencies between the unit treatments are bounded, then $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are sums of variables whose dependencies will be similarly bounded. For this reason, under a variety of designs we may expect the values of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$, while unknown due to $\mathcal{I}$} being unknown, to concentrate at their expectation (which equals $\tau^{\textup{basic}}$) and to be asymptotically normal.
Our motivation for $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ is the following: under mild conditions on the experiment design, the maximum of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ is asymptotically lower bounded by the magnitude of the propensity-weighted difference in outcomes between treated and control, given by
$\Delta=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i},$
as stated by Proposition 1 below:
Proposition 1.
@@
Let the total weights of the treated and control converge to their expectations, so that
$\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}=N+O_{P}(N^{1/2})\quad\textup{and}\quad\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}=N+O_{P}(N^{1/2})$ (6)
Then it holds that
$\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1},\hat{\tau}_{2})+O_{P}(N^{1/2})$ (7)
As a corollary, it immediately follows that if $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are consistent for $\tau^{\textup{basic}}$, then $\max(\hat{\tau}_{1},\hat{\tau}_{2})$ will be consistent as well, and $|\Delta|$ will be an asymptotic lower bound for $\tau^{\textup{basic}}$.
Corollary 1.
Let the conditions of Theorem 1 hold. If $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ are within $O_{P}(N^{1/2})$ of $\tau^{\textup{basic}}$, then
$\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|\leq\tau^{\textup{basic}}+O_{P}(N^{1/2})$ (8)
We note that $\Delta$, which involves units outside of $\mathcal{I}$}, need not concentrate in order for (7) or (8) to hold. In contrast to $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$, which depend only on units in $\mathcal{I}$}, it may well be that $\Delta$ becomes less stable as the number of participants grows, but nonetheless is asymptotically a valid lower bound on the estimand.
Discussion
To give intuition for Theorem 1, consider an experiment in which $100$ units are randomly divided into equal-sized groups for treatment and control, with average outcomes of $0.4$ (20/50) for the treatment group and $0.8$ (40/50) for the control. How large can $|{\mathcal{I}}|$ be? Units in $\mathcal{I}$} have constant outcomes and are equally likely to be assigned to treatment or control. It follows that up to sampling variation, such units should be evenly distributed between treatment and control for each outcome level. Since there are 20 treated units and 40 control units with outcome level 1, it trivially holds that at most all 20 of the treated units with outcome level 1 could belong to $\mathcal{I}$}, and thus at most roughly 20 out of the 40 control units with outcome level 1 could belong to $\mathcal{I}$} as well. (Otherwise, units in $\mathcal{I}$} with outcome level 1 would not be evenly distributed between treatment and control.) Likewise, since only 10 control units have outcome level 0, at most all 10 could belong to $\mathcal{I}$}, implying that at most roughly 10 out of the 30 treated units with outcome level 0 could belong to $\mathcal{I}$} as well. This yields an upper bound of $(20+10)\times 2=60$ units who are unaffected by treatment, or equivalently at least 40 units who are affected by treatment. This lower bound of 40 equals the difference in outcomes of 0.4 between treatment and control, and also equals the maximum possible value of $\hat{\tau}_{1}$ over all hypotheses for the unknown $\mathcal{I}$}. The limiting quantities that were used to construct this bound, namely the 20 treated units with outcome equal to 1, and the 10 control units with outcome equal to 0, correspond to the units $\{i:H_{i}=1\}$.
When might this lower bound on $\tau$ be tight? Can $\tau$ be upper bounded? To build intuition, we continue the example by presenting a generative model under which the lower bound is tight. In this model, the contrast variable $X$ appearing in (7) is a perfect predictor of the outcome for units not in $\mathcal{I}$. We then describe two models under which the lower bound is loose, and which suggest that non-vacuous upper bounds on $\tau$ cannot be constructed without additional assumptions.